OpenStudy (anonymous):

Use a lemon to prove that if 7x + 4 is even, then 3x - 11 is odd and X C- Z.

OpenStudy (anonymous):

@ganeshie8 @myininaya @UnkleRhaukus @RadEn @robtobey @chmvijay @Luigi0210 @eliassaab @genius12 @Vincent-Lyon.Fr @Preetha @Kainui @ehuman @wolfe8 @lucaz @INeedHelpPlease? @khadeeja @thadyoung @A_clan @tester97 @Andras @Confusionist @linh412986 @BlackLabel @leozap1 @JMark @Microrobot @zacharyf @Rubio101 @NaomiBell1997 @MeganEdward @theballer225 @link,zoro @Goldenmoon @CrayolaCrayon_

OpenStudy (anonymous):

\[\bf 7x+4 \equiv 0 (\mod 2) \implies 7x \equiv 0 (\mod 2) \implies x \equiv 0 (\mod 2)\]Hence 'x' is always even in order for 7x + 4 to be even. Now let \( x = 2k, for \ k \ge 0\). Then: \[\bf 3x-11=3(2k)-11=2(3k)-2(5)-1=2(3k-5)-1\]Therefore 3x - 11 will always be odd when 7x + 4 is even. Q.E.D. @FutureMathProfessor

OpenStudy (anonymous):

gn for now. bye gais.

OpenStudy (anonymous):

Night! I suppose I better go to bed too because I hate sleeping when it's not black outside LOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOL

OpenStudy (kc_kennylau):

For the last time, stop using mass tags.

OpenStudy (kc_kennylau):

Tips: Use \ ( \in \ ) to make \(\in\) :)

OpenStudy (anonymous):

/in

OpenStudy (anonymous):

\in

OpenStudy (anonymous):

( \in \ )

OpenStudy (anonymous):

How @kc_kennylau

OpenStudy (anonymous):

\in \

OpenStudy (anonymous):

\ ( \in \ )

OpenStudy (anonymous):

( in \ )

OpenStudy (kc_kennylau):

Remove all the spaces in \ ( \in \ )

OpenStudy (kc_kennylau):

Because If I did so it wouldn't give u the code

OpenStudy (anonymous):

\(\in\)

OpenStudy (anonymous):

WHOA.

OpenStudy (kc_kennylau):

exactly

OpenStudy (anonymous):

\(\in\)\(\in\)\(\in\)\(\in\)\(\in\)\(\in\)\(\in\)\(\in\)\(\in\)\(\in\)\(\in\)\(\in\)\(\in\)\(\in\)\(\in\)\(\in\)\(\in\)\(\in\)\(\in\)\(\in\)\(\in\)\(\in\)\(\in\)\(\in\)\(\in\)\(\in\)\(\in\)\(\in\)\(\in\)\(\in\)\(\in\)\(\in\)\(\in\)\(\in\)\(\in\)\(\in\)\(\in\)\(\in\)\(\in\)\(\in\)\(\in\)\(\in\)\(\in\)\(\in\)\(\in\)\(\in\)\(\in\)\(\in\)\(\in\)\(\in\)\(\in\)\(\in\)\(\in\)\(\in\)\(\in\)\(\in\)\(\in\)\(\in\)\(\in\)\(\in\)\(\in\)\(\in\)\(\in\)\(\in\)\(\in\)\(\in\)\(\in\)\(\in\)\(\in\)\(\in\)\(\in\)\(\in\)\(\in\)\(\in\)\(\in\)

OpenStudy (kc_kennylau):

bye i have to go

OpenStudy (anonymous):

Show me more stuff

OpenStudy (anonymous):

\(\subseteq\)

OpenStudy (kc_kennylau):

really have to go have fun

OpenStudy (anonymous):

Wow! This is like learning how to thrust your wife on your wedding night! Im gona have fun with this!